# Eccentric features in the gravitational wave phase of dynamically formed black hole binaries

### v1

@@ 3.3 Generalised Behaviour
We generalise our results more by showing the transition between moderate and high eccentricities and when strong curvature effects begin to occur. In the top panel of Fig. 10 we show $d\phi_{\mathrm{max}}$ as a function of $f_{m}$ for a range of $e_{\mathrm{out}}$ (from $0.6$ to $0.99$). Each horizontal slice of this plot can be viewed as a curve like Fig. 6. It becomes clear that eccentricity effects are already visible as low as $e_{\mathrm{out}}=0.6$. There, the location of the maximum phase shift has in fact already been slightly shifted to the left of periastron. Between $e_{\mathrm{out}}=0.6$ and $e_{\mathrm{out}}=0.95$, the maximum possible phase shift steadily increases by about 3 orders-of-magnitude.
It is further informative to show the emergence of the secondary peak as we increase the outer eccentricity, and in what range of $f_{m}$ the peak appears. We portray this in the bottom panel of Fig. 10; the $x$-axis depicts $f_{m}$ and the $y$-axis $e_{\mathrm{out}}$. We define the presence of a double peak by a sign change in the second time derivative of the phase shift. The green area shows the region of $f_{m}$ and $e_{\mathrm{out}}$ for which there is a secondary peak. The double peak does not occur for the lowest eccentricities, as the curvature effects of passing pericentre are not large enough. With this setup, the secondary peak appears for $e_{\mathrm{out}}$ as low as $\sim 0.7$. As we increase $e_{\mathrm{out}}$ the green region widens. On the left, it asymptotes towards apocentre. On the right, the width increases faster due to the strong curvature effects. Above $e_{\mathrm{out}}\sim 0.9$, the width decreases again, as the binary moves extremely fast close to pericentre. For an eccentricity of $\sim 0.95$, the time between pericentre passage and a merger happening at $f_{m}$ of $\sim-0.5\pi$ is so short that the two peaks become indistinguishable.
In summary, the features of the phase shift that are typical of triples with a high outer eccentricity already show up in the phase shift at moderate to moderately high eccentricity. In terms of detections this is a positive result as it indicates towards the possibility of distinguishing circular outer orbits from their eccentric counterparts not just in the most extremely eccentric cases. Below, we will illustrate the use of this model by applying it to case studies in realistic astrophysical environments.
@@ 4 Astrophysically Relevant Scenarios
Chaotic scatterings are ideal environments for producing systems with potentially detectable phase shifts. In the top panel of Fig. 11, we show once again the trajectories of a chaotic triple producing a highly eccentric merger. From the positions and velocities at merger we can extract the orbital elements of the triple that are needed as input for our model to compute the phase shift, except for $a_{0}$ and $e_{0}$ which we extract at some reference time before merger. In the bottom panel of Fig. 11 we plot the binary and reference trajectory (blue and orange) on top of the data from the N-body simulation (grey shaded).
4.1 Triple Scattering
Figure 11: Top: The trajectories of the scattering from Fig. 1 for illustration. Bottom: the binary and reference trajectories including Rømer delay for this particular scattering, including the extracted ellipse of the Keplerian outer orbit (dashed). The final orbit of the binary COM and third BH generated in the N-body scattering are depicted as shaded lines. The location of the maximum phase shift is depicted by the red cross. All trajectories are shown in the COM frame of the triple.
Figure 12: Phase shift as a function of time (top) and orbital eccentricity (bottom) for the case study depicted in Fig. 11. In blue, we show the result from our semi-analytical model, and in pink our approximations: the solid line depicts the circular approximation (Eq. 23), and the dashed line shows the fast linear approximation (Eq. 29).
The resulting phase shift of this scattering is depicted in Fig. 12, as a function of time (top) and orbital eccentricity (bottom). We observe that both versions of our approximation slightly overestimate the magnitude of the (maximum) phase shift, as compared to the prediction from our full semi-analytical model. The reason for this is simply that for this setup and in this part of the outer orbit, the acceleration when the phase shift is at its maximum (red cross) has changed significantly compared to the value at merger. However, despite these known limitation of our analytical approximations, we do find that the estimates for $d\phi_{\mathrm{max}}$ from both the fast linear approximation (Eq. 32) and the circular approximation (Eq. 23) are only about a factor of $1-3$ off compared to our more accurate semi-analytical model. This motivates greatly to explore the possibilities of using such approximations for future large-data-projects involving determining GW phase shifts for millions of PN few-body scatterings across cosmic time for e.g. different cluster environments. Lastly, we see that our estimates here are fully consistent with the numerically estimated phase shift for this scattering shown in Samsing et al. (2024). Especially our semi-analytical model correctly captures the slope of $d\phi$ on both sides of the peak, whereas the circular approximation that was over-plotted in Samsing et al. (2024) deviates from the truth at earlier times.
@@ 4.2 Influence of Tides
BBHs assembled during chaotic scatterings are often undergoing their first part of their inspiral while being subject to tidal influence from the nearby perturber. This effect was studied for a few numerical cases in Samsing et al. (2024). Here we perform a more controlled study of this, by investigating how the tidal influence on the eccentricity of the inspiralling BBH will propagate to changes in the expected later GW phase shift. As described below, for this we consider a BBH that is assembled near the Hill sphere with respect to the perturber.
For our initial setup, we consider two BHs that become bound after radiating an amount $\Delta E_{\text{GW}}$ of GW radiation during their first pericentre passage. We can describe this as (see Peters, 1964; Hansen, 1972; Samsing et al., 2018):
$\Delta E_{\text{GW}}=\frac{85\pi}{12\sqrt{2}}\frac{G^{7/2}}{c^{5}}\frac{m_{1}^{2}m_{2}^{2}m^{1/2}}{r_{p}^{7/2}},$ (34)
Here, $r_{p}=a(1-e)$ is the distance at pericentre, and $m=m_{1}+m_{2}$. Under the assumption that the two BHs will end up near the Hills sphere after their first pericenter passage, they will have an orbital energy given by,
$\begin{split}\Delta E_{\text{orb}}&=\frac{Gm^{2}}{2a}\approx\frac{Gm^{2}}{R_{H}}.\end{split}$ (35)
The definition of the Hill radius is
$\begin{split}R_{H}&=r(f)\left(\frac{m_{1}}{3m_{3}}\right)^{1/3}\\ &=\frac{a_{\text{out}}(1-e^{2}_{\text{out}})}{1+e_{\text{out}}\cos f}\left(\frac{m_{1}}{3m_{3}}\right)^{1/3}.\end{split}$ (36)
By combining Eqs. 34 and 35 we now have an estimate for the initial eccentricity and semi-major axis given the BBH forms through a capture near the Hill sphere,
$e_{0}\approx 1-2\left(\frac{85\pi}{12\sqrt{2}}\right)^{2/7}\frac{G^{5/7}}{c^{10/7}}\frac{m_{1}^{4/7}m_{2}^{4/7}}{m^{3/7}}\frac{1}{R_{H}^{5/7}}.$ (37)
@@
Figure 13: Phase shift as a function of time (top) and peak GW frequency (bottom) for four different initial binary eccentricities. We start with $e_{0}$ according to Eq. 37, and systematically subtract a small value to alter the phase shift evolution. We use $M_{1}=M_{2}=5M_{\odot}$, $M_{3}=20M_{\odot}$, $e_{\mathrm{out}}=0.9$ and $a_{\mathrm{out}}=10R_{\odot}$.
Upon formation of the inspiralling BBH, when the binary is still relatively wide, tidal forces from the third object can alter the initial semi-major axis and eccentricity (see also Samsing et al., 2024). Exactly how and to what extent the binary is influenced by the tides is chaotic and is dependent on several factors, such as the orientation at formation with respect to the perturber or distance to the perturber. Any change to the initial conditions may have a large effect on the inspiral time and thereby evolution of the phase shift. To illustrate this, here we investigate how small changes in the initial binary eccentricity $e_{0}$ may affect the phase shift evolution. We create a binary that forms on the Hill radius with eccentricity and semi-major axis according to Eqs. 37 & 38. The binary forms at a true anomaly $f_{f}=0.1\pi$. We evolve this system 4 times; once with the original $e_{0}$ and $a_{0}$, and thrice with a slight alteration of the initial eccentricity that can arise from the tidal forces of the perturber. We decrease the initial eccentricity by $10^{-5}$, $10^{-4}$ and $10^{-3}$. The trajectories of each of these four systems are depicted in Fig. 14. In Fig. 13, we show their phase shift evolution as a function of time and peak GW frequency.
A small change in eccentricity may lead to drastic changes in the initial pericentre distance, especially in mergers that form at such high eccentricities. The system that we study here has an initial eccentricity of $e_{0}\sim 0.998$, and decreasing this by only $10^{-3}$ would increase the pericentre distance by a factor of 1.6. This allows for less power emitted in GWs in these early stages, therefore resulting in a longer inspiral time. With an eccentricity change of $10^{-4}$, the time to coalescence becomes $\sim 1.2$ times larger. For $10^{-3}$, we obtain an increase as large as a factor $\sim$4.6. In terms of detectability, this might work in our favour. However, as we see in the bottom panel of Fig. 13, this also has its downsides as it shifts the evolution of the phase shift to a lower peak GW frequency, potentially pushing it out of band. Given that
@@
increasing $r_{p}$ by a factor 1.6 lowers the peak GW frequency at formation by a factor $\sim$2.05.
In conclusion, phase shifts may get largely amplified by tides from the perturber tidal effects, potentially increasing chances of a detection. Small drops in initial eccentricity lead to longer inspiral times and higher phase shifts as well as a shift to lower peak frequencies. The latter might result in the binary being pushed out of band for a LIGO-like detector.
Figure 14: Trajectories of the binary COM (turquoise), reference binary (orange), and third BH (grey), for 3 different values subtracted from $e_{0}$ ($10^{-5}$, $10^{-4}$, and $10^{-3}$). The choices for the other parameters are described in Sec. 4.2.
5 Summary & future outlook
In this paper, we developed for the first time a model for the Rømer delay-induced phase shift of an eccentric BBH that is on an eccentric orbit around a perturber. This is an extension of Samsing et al. (2024) in which we made the first analytical model for phase shifts of eccentric BBHs on circular outer orbits. Next to the full calculation, we also presented several analytical approximations. With these tools, we investigated in detail the effects of the phase shift that are unique to the outer orbit being eccentric. Lastly, we applied our model to relevant astrophysical scenarios of chaotic 3-body scatterings and studied the effect of tides from the third-body onto the phase shift.
@@ 5 Summary & future outlook
  • For BBHs that inspiral along eccentric orbits the time evolution of the phase shift is dependent on where in the orbit the binary evolves. For low and moderate outer eccentricities, the leading order effect is the amplitude of the peak phase shift which $\propto r_{m}^{-2}$, which shows that the phase shift is highest when the binary merges at pericentre, and lowest when it merges at apocentre.
@@
  • At high outer eccentricities ($e_{\mathrm{out}}=0.9$ in our example), effects of the strong curvature around pericentre are imprinted onto the phase shift. Two characteristics stand out: first, the maximum possible phase shift no longer occurs when the binary merges at pericentre, but slightly beyond ($f_{m}\sim-0.2\pi)$. Secondly, in certain parts of parameter space ($f_{m}\sim-0.75\pi$ in our example), two peaks may occur. If the binary passes pericentre before the main peak happens, a boost in phase shift due to the large curvature and high velocity may give rise to a secondary peak. These features are important properties of eccentric outer orbits and may arise at outer eccentricities as low as $e\sim 0.7$ with our setup.
@@
  • Chaotic scatterings are ideal environments for such phase shifts and their orbits have to be described by our eccentric models, compared to the standard circular.
@@
  • In the case of GW capture formatioin of BBHs at the Hill sphere, tides from the third body may give rise to small changes in the initial binary eccentricity. This induces an increased time to coalescence and thereby a higher maximum $d\phi$. At the same time, this pushes the peak frequency down, potentially out of band.
The completion of this work has opened up the window to numerous interesting future studies that we intend to carry out or are in the process of working on. We list the most important ones:
  • Quantifying the detectability. This is a very important piece in showcasing the relevance of studying this effect in more detail. A vital next step is to investigate extensively in what scenarios this phase shift might be observable in GW signals. Full parameter estimations should be executed with state-of-the art eccentric waveforms to make quantitative claims about this. This includes making comparisons between different (generations of) GW detectors, identifying different ranges of values the parameters can take depending on the formation environment, and possible degeneracies between parameters.
  • Phase shifts from chaotic scatterings. Our case study in Sec. 4.1 is only an example of the many triple scatterings that occur in stellar clusters, leading to dynamically formed mergers. Using full numerical codes of GCs such as CMC can allow us to make claims of populations of triples and how their properties are imprinted in the phase shifts they produce. This is an important puzzle piece in gaining a better understanding of the formation environments of GW sources.
  • Hyperbolic scatterings. Environments such as the above may produce many 3-body scatterings where the outer orbit is hyperbolic. Extending our model to the hyperbolic case will pave the way to also study phase shifts from these interactions, including detectability and rates.
  • Other types of environments. Clusters with chaotic scatterings are not the only environment where detectable Rømer phase shifts might occur. Other environments, such as hierarchical triples or AGN disks with a central supermassive BH may also produce the effect we studied in this paper. This may give rise to different features and properties that are unique to those specific sub-channels. As such, more detailed studies on those environments can potentially provide valuable results.
  • Other types of phase shifts. Naturally, Rømer delay is not the only source of dephasing for (eccentric) GW sources. Numerous other effects can leave imprints onto the waveform: Shapiro delay, gravitational redshift, Kozai-Lidov effects, gravitational lensing, et cetera. It will be very important to list each of these types, and understand their characteristics, relative importance in different binary formation scenarios and degeneracies.
  • Observer dependence. In each of the above points, it is assumed that the observer dependence will be built into our phase shift model. Here we mention that it is useful to also separately research the dependence of the phase shift on the observer position, and reveal potential observation-related challenges that might arise.
6 Acknowledgements

### v2

@@ 3.3 Generalised Behaviour
We generalise our results more by showing the transition between moderate and high eccentricities and when strong curvature effects begin to occur. In the top panel of Fig. 10 we show $d\phi_{\mathrm{max}}$ as a function of $f_{m}$ for a range of $e_{\mathrm{out}}$ (from $0.6$ to $0.99$) for our considered system. Each horizontal slice of this plot can be viewed as a curve like Fig. 6. It becomes clear that eccentricity effects are already visible as low as $e_{\mathrm{out}}=0.6$. There, the location of the maximum phase is already shifted to the left of periastron. Between $e_{\mathrm{out}}=0.6$ and $e_{\mathrm{out}}=0.95$, the maximum possible phase shift steadily increases by about 3 orders-of-magnitude.
It is further informative to show the emergence of the secondary peak as we increase the outer eccentricity, and in what range of $f_{m}$ the peak appears. We portray this in the bottom panel of Fig. 10, where the $x$-axis shows $f_{m}$ and the $y$-axis $e_{\mathrm{out}}$. We define the presence of a double peak by a sign change in the second time derivative of the phase shift. The green area shows the region of $f_{m}$ and $e_{\mathrm{out}}$ for which there is a secondary peak. The double peak does not occur for the lowest eccentricities, as the curvature effects of passing pericentre are not large enough. With this setup, the secondary peak appears for $e_{\mathrm{out}}$ as low as $\sim 0.7$. As $e_{\mathrm{out}}$ increases the green region widens. On the left, it asymptotes towards apocentre. On the right, the width increases faster due to the strong curvature effects. Above $e_{\mathrm{out}}\sim 0.9$, the width decreases again, as the binary moves extremely fast close to pericentre. For an eccentricity of $\sim 0.95$, the time between pericentre passage and a merger happening at $f_{m}$ of $\sim-0.5\pi$ is so short that the two peaks become indistinguishable.
In summary, the features of the phase shift that are typical of triples with a high outer eccentricity can show up in the phase shift at moderate to moderately high eccentricity. In terms of detections this is a positive result as it indicates towards the possibility of distinguishing circular outer orbits from their eccentric counterparts not just in the most extremely eccentric cases.
@@ 4 Astrophysically Relevant Scenarios
4.1 Three-body Scatterings
Chaotic scatterings are ideal environments for producing systems with potentially detectable GW phase shifts. In the top panel of Fig. 11, we show once again the trajectories of a chaotic triple producing a highly eccentric merger. From the positions and velocities at merger we can extract the orbital elements of the triple that are needed as input for our model to compute the phase shift, except for $a_{0}$ and $e_{0}$ which we extract at some reference time before merger. In the bottom panel of Fig. 11 we plot the binary and reference trajectory (blue and orange) on top of the data from the N-body simulation (grey shaded).
The resulting phase shift $d\phi$ of this scattering is depicted in Fig. 12, as a function of time (top) and orbital eccentricity (bottom). We observe that both versions of our approximation slightly overestimate the magnitude of the (maximum) phase shift, as compared to the prediction from our full semi-analytical model. The reason for this is simply that for this setup, and in this part of the outer orbit, the acceleration when the phase shift is at its maximum (red cross) has changed significantly compared to the value at merger. However, despite these known limitation of our analytical approximations, we do find that the estimates for $d\phi_{\mathrm{max}}$ from both the fast linear approximation (Eq. 31) and the circular approximation (Eq. 22) are only about a factor of $1-3$ off compared to our more accurate semi-analytical model. This greatly motivates to explore the possibilities of using such approximations for future large-data projects involving determining GW phase shifts for millions of PN few-body scatterings across cosmic time for e.g. different cluster environments. Lastly, we see that our estimates here are fully consistent with the numerically estimated phase shift for this scattering shown in Samsing et al. (2024). Especially our semi-analytical model correctly captures the slope of $d\phi$ on both sides of the peak, whereas the circular approximation that was over-plotted in Samsing et al. (2024) deviates from the truth at earlier times.
Figure 11: Eccentric orbits in 3-body interactions. Top: The trajectories of the scattering from Fig. 1 for illustration. Bottom: the binary and reference trajectories including Rømer delay for this particular scattering, including the orbit for our circular approximation. The shaded dashed lines show the final orbits of the binary (blue) and third object (green) in the N-body simulation. The location of the maximum phase shift is depicted by the red cross. All trajectories are shown in the COM frame of the triple.
Figure 12: 3-body GW phase shifts and models. GW phase shift as a function of time (top) and orbital eccentricity (bottom) for the case study depicted in Fig. 11. In blue, we show the result from our semi-analytical model, and in pink our approximations: the solid line depicts the circular approximation (Eq. 22), and the dashed line shows the fast linear approximation (Eq. 28).
@@ 4.2 Influence of Tides
BBHs assembled during chaotic scatterings are often undergoing their first part of their inspiral while being subject to tidal influence from the nearby perturber. This effect was studied for a few numerical cases in Samsing et al. (2024). Here we perform a more controlled study of this, by investigating how the tidal influence on the eccentricity of the inspiralling BBH will propagate to changes in the expected later GW phase shift. As described below, for this we consider a BBH (BH1,BH2) that is assembled near the Hill sphere with respect to a third object (BH3).
For our initial setup, we consider two BHs that become bound near BH3 after radiating an amount $\Delta E_{\text{GW}}$ of GW radiation during their first pericentre passage. We can describe this as (see Peters, 1964; Hansen, 1972; Samsing et al., 2018):
$\Delta E_{\text{GW}}=\frac{85\pi}{12\sqrt{2}}\frac{G^{7/2}}{c^{5}}\frac{m_{1}^{2}m_{2}^{2}m_{12}^{1/2}}{r_{p}^{7/2}},$ (33)
Here, $r_{p}=a(1-e)$ is the distance at pericentre, and $m_{12}=m_{1}+m_{2}$. Under the assumption that the two BHs will end up near the Hill sphere relative to BH3 after their first pericentre passage, they will have an orbital energy given by,
$\begin{split}\Delta E_{\text{orb}}&=\frac{Gm_{1}m_{2}}{2a}\approx\frac{Gm_{1}m_{2}}{R_{H}}.\end{split}$ (34)
The definition of the Hill radius of BH1, on which we place BH2, is
$\begin{split}R_{H}&=r(f)\left(\frac{m_{1}+m_{2}}{3m_{3}}\right)^{1/3}\\ &=\frac{a_{\text{out}}(1-e^{2}_{\text{out}})}{1+e_{\text{out}}\cos f}\left(\frac{m_{1}+m_{2}}{3m_{3}}\right)^{1/3}.\end{split}$ (35)
By combining Eqs. 33 and 34 we now have an estimate for the initial eccentricity and semi-major axis in this GW capture scenario, which is a toy description of what do realistically happen in triple scatterings,
$e_{0}\approx 1-2\left(\frac{85\pi}{12\sqrt{2}}\right)^{2/7}\frac{G^{5/7}}{c^{10/7}}m_{1}^{2/7}m_{2}^{2/7}m_{12}^{1/7}\frac{1}{R_{H}^{5/7}},$ (36)
@@
Figure 13: Orbital changes from tides. Trajectories of the binary COM (turquoise), reference binary (orange), and third BH (grey), for 2 different values subtracted from $e_{0}$ ($10^{-4}$, and $10^{-3}$). The choices for the other parameters are described in Sec. 4.2.
Figure 14: Tidal influence on GW phase shift evolutions. Phase shift as a function of time (top) and peak GW frequency (bottom) for three different initial binary eccentricities. We start with $e_{0}$ according to Eq. 36, and systematically subtract a small value to alter the phase shift evolution. We use $m_{1}=m_{2}=5M_{\odot}$, $m_{3}=20M_{\odot}$, $e_{\mathrm{out}}=0.9$ and $a_{\mathrm{out}}=10R_{\odot}$. We observe that each of these trajectories shows the secondary peak in the phase shift that was discussed in the above.
Upon formation of the inspiralling BBH, when the binary is still relatively wide, tidal forces from the third object can alter the initial semi-major axis and eccentricity (see also Samsing et al., 2024). Exactly how and to what extent the binary is influenced by the tides is chaotic and is dependent on several factors, such as the orientation and distance with respect to the third object. Any change to the initial conditions may have a large effect on the inspiral time and thereby evolution of the phase shift. To illustrate this, here we investigate how small changes in the initial binary eccentricity $e_{0}$ may affect the phase shift evolution. We create a binary that forms on the Hill radius with eccentricity and semi-major axis according to Eqs. 36 and 37. The binary forms at a true anomaly $f_{f}=0.1\pi$. We evolve this system 3 times: once with the original $e_{0}$ and $a_{0}$, and twice with a slight alteration of the initial eccentricity that can arise from the tidal forces of the perturber. We decrease the initial eccentricity by $10^{-4}$ and $10^{-3}$. The trajectories of each of these four systems are depicted in Fig. 13. In Fig. 14, we show their phase shift evolution as a function of time and peak GW frequency.
As seen, a small change in eccentricity may lead to drastic changes in the initial pericentre distance, especially in mergers that form at such high eccentricities. The system that we study here has an initial eccentricity of $e_{0}\sim 0.998$, and decreasing this by only $10^{-3}$ would increase the pericentre distance by a factor of 1.49. This allows for less power emitted in GWs in these early stages, therefore resulting in a longer inspiral time. With an eccentricity change of $10^{-4}$, the time to coalescence becomes $\sim 1.17$ times larger. For the $10^{-3}$ case we obtain an increase in time as large as a factor $\sim$3.7. In terms of detectability, this might work in our favour. However, as we see in the bottom panel of Fig. 14, this also has its downsides as it shifts the evolution of the phase shift to a lower peak GW frequency, potentially pushing it out of band. Given that
@@
increasing $r_{p}$ by a factor 1.49 lowers the peak GW frequency at formation by a factor $\sim$1.8.
In conclusion, phase shifts may get largely amplified by tides from the perturber tidal effects, potentially increasing chances of a detection. Small drops in initial eccentricity lead to longer inspiral times and higher phase shifts as well as a shift to lower peak frequencies. The latter might result in the binary being pushed out of the detector band.
5 Summary
In this paper, we explored for the first time the Rømer delay-induced phase shift, arising from an acceleration to the binary COM, of an eccentric BBH that is on an eccentric orbit around a third object. Besides presenting results from a semi-analytical model including orbital evolution using PN-orbit averaged equations and Kepler’s Equations, we further presented several analytical approximations. With these tools, we investigated the effects of the GW phase shift that are unique to the outer orbit being eccentric together with an eccentric evolving inner orbit merging BBH. Lastly, we applied our model to relevant astrophysical scenarios of chaotic 3-body scatterings, and studied the effect of tides from the third-body onto the GW phase shift.
@@ 5 Summary
  • For BBHs that inspiral along eccentric orbits the time evolution of the phase shift is dependent on where in the orbit the binary evolves. For low and moderate outer eccentricities, the leading order effect relates to how the magnitude of the GW phase shift can change over the orbit with the distance between the binary and third object, $r_{m}$, as $\propto r_{m}^{-2}$. This especially illustrates that the GW phase shift easily can change by order-of-magnitude over the orbit with its maximum around merger at pericentre.
@@
  • At high outer eccentricities ($e_{\mathrm{out}}=0.9$ in our example), effects of the strong curvature around pericentre are imprinted onto the phase shift. Two characteristics stand out: first, the maximum possible phase shift no longer occurs when the binary merges at pericentre, but slightly beyond ($f_{m}\sim-0.2\pi)$. Secondly, in certain parts of parameter space ($f_{m}\sim-0.75\pi$ in our example), two peaks may occur. If the binary passes pericentre before the main peak happens, a boost in phase shift due to the large curvature and high velocity may give rise to a secondary peak. These features are important properties of eccentric outer orbits..
@@
  • Chaotic scatterings are ideal environments for such phase shifts and their orbits have to be described by eccentric models, as the ones we presented here, compared to the standard circular.
@@
  • In the case of GW capture formation of BBHs at the Hill sphere, tides from the third body may give rise to small changes in the initial binary eccentricity. This induces an increased time to coalescence and thereby a higher maximum $d\phi$. At the same time, this pushes the peak frequency down, potentially out of band.
The completion of this work has opened up the window to numerous interesting future studies, including quantifying the detectability, GW phase statistics from cluster simulations, comparisons to other dynamical environments, as well as incorporating observer dependence and other GR GW phase shift effects. We reserve that for future works.
Acknowledgements